OpenStudy (anonymous):

f(x)=610x^2-10,250x+1,235 determine the max or min and write answer as a coordinate????

OpenStudy (mertsj):

I see it has a minimum. Do you see that?

OpenStudy (anonymous):

yes because it's postive right

OpenStudy (mertsj):

Yes.

OpenStudy (mertsj):

And that minimum is the vertex. Right?

OpenStudy (anonymous):

right

OpenStudy (mertsj):

And the x coordinate of the vertex can easily be found because we know that x = -b/2a Right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

oh by the way it says x=.28 and then it says to determine the min and write as coordinate, idk if .28 makes any difference

OpenStudy (mertsj):

And -b/2a= 10250/1220=8.40

OpenStudy (mertsj):

Maybe you'd better post the entire problem.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

evaluate each expression below for x=.28. determine the max or min of f(X)=610x^2-10250x+1235. write answer as a coordinate.

OpenStudy (mertsj):

Ok. We can do that at the end. Right now we're in the middle of finding the vertex. So do you agree that the x coordinate of the vertex is 8.40?

OpenStudy (anonymous):

yes i agree

OpenStudy (mertsj):

So now we can find the y coordinate of the vertex by replacing x with 8.40 and evaluating the expression.

OpenStudy (mertsj):

I got -41823.4 What did you get?

OpenStudy (anonymous):

-4.18234

OpenStudy (mertsj):

How did you get a decimal point after the 4?

OpenStudy (anonymous):

let me calculate it again ok

OpenStudy (mertsj):

The numbers are right. just the decimal is off.

OpenStudy (mertsj):

maybe your calculator is in scientific notation mode and it has an exponent of 4. Could that be?

OpenStudy (anonymous):

is cuz i set it to round and it's literallly going crazy so let's go with ur answer ok

OpenStudy (mertsj):

Ok. Now we know the vertex is (8.4,-41823.4)

OpenStudy (mertsj):

And now we have to plug in the .28 and find the value of the expression for that x value.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

I got -6917.176 What did you get?

OpenStudy (anonymous):

-6.917176

OpenStudy (anonymous):

the decimal again

OpenStudy (mertsj):

you've got to get your calculator out of scientific notation mode. Do you have a EE button?

OpenStudy (mertsj):

It automatically gives you your answers in scientific notation and you are overlooking the number at the upper right of the screen that tells you the exponent of the 10

OpenStudy (anonymous):

i redid it and now i got -1.59??????

OpenStudy (mertsj):

I'll do it again. I'll bet that every problem you do, you get an answer between 1 and 10. Am I right? Enter 20 times 500

OpenStudy (anonymous):

1.00

OpenStudy (mertsj):

There you go. Of course 20 times 500 is 10000 not 1. Did you look for a little number at the upper right of your screen? It should be a 4

OpenStudy (mertsj):

In any case, I've double checked it a number of times. f(.28)=-1587.176

OpenStudy (anonymous):

so ok then how do i write it as a coordinate?

OpenStudy (mertsj):

(.28,-1587.176)

OpenStudy (mertsj):

And for the vertex: Ok. Now we know the vertex is (8.4,-41823.4)

OpenStudy (anonymous):

yes ok i'm clear now

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

thank you!!!!!!!!

OpenStudy (mertsj):

yw